# Understanding Discrepancies of Wavefunction Theories for Large Molecules

### v1

@@ 2.3 Estimating the overcorrelation of (T) for weak interactions
In summary, Table 1 and Figs.2 & 3 demonstrate that CCSD(cT) theory achieves excellent agreement for noncovalent interaction energies between molecular complexes compared to DMC and CCSDT theory. However, we stress that the CC series of methods (CCSD, CCSDT and CCSDTQ) is observed to yield monotonic and rapidly converging interaction energies for small and weakly bound complexes [37]. Based on this knowledge, we emphasize that the Q contribution to the interaction energies can be expected to be smaller than its T counterpart, but could possibly yield a significant contribution. Indeed, this is part of the reason for the success of the CCSD(T) approximation for very small molecules, where CCSD(T) is often fortuitously closer to CCSDTQ than CCSDT [37]. Here, we argue that this error cancellation no longer functions in the case of large molecular complexes involving strongly polarizable systems such as C2C2PD, C3GC and C60@[6]CPPA. A similar problem is known to occur in MP2 theory, where the truncation of the perturbation series also leads to significantly too strong interaction energies for systems with large polarizability, although MP2 yields relatively accurate interaction energies for systems with an intermediate polarizability [31].
@@
Having demonstrated and explained the reasons for the overestimation of absolute interaction energies on the level of CCSD(T) theory for small molecules with up to 24 atoms and the C2C2PD system, we now want to turn to the discussion of the remaining large molecular complexes where significant absolute discrepancies between DMC and CCSD(T) have been observed. These systems include C3GC from the L7 data set and the C60@[6]CPPA buckyball-ring. Here, substantial differences in the binding energies of and were reported, respectively after subtracting error bars. Although CCSD(cT) calculations for systems of that size are currently not feasible using our approach, we now introduce a simplified model that allows us to estimate the change in interaction energies from CCSD(T) to CCSD(cT) in an approximate manner. Given the linear trend between the different correlation energy contributions to the interaction energy depicted in Fig. 4, it is possible to estimate the (cT) contribution for systems where only MP2, CCSD and (T) are known. These numbers can be calculated using a computationally efficient LNO-CCSD(T) implementation [40]. Results computed in this manner are denoted as CCSD(cT)-fit. Details about this procedure and the corresponding error estimates are provided in the supplementary information.
| System | CCSD(T) | LNO-CCSD(T) [10] | CCSD(cT) | CCSD(cT)-fit | DMC [10] | DMC [11] |
@@
Table 2 gives our estimated CCSD(cT) interaction energies in comparison to CCSD(T) and DMC findings for seven large molecular complexes. A comparison between CCSD(cT)-fit and the explicitly calculated CCSD(cT) results for GGG, GCGC and C2C2PD shows that the linear regression model is sufficiently reliable for the systems studied in this work. For comparison Table 2 also summarizes the DMC interaction energies from Refs. [11] and [10], which agree to within at least for GGG, C2C2PD, PHE and C3GC. For the remaining systems the DMC estimates show a larger discrepancy and for C60@[6]CPPA only one DMC estimate is available. Although the DMC binding energies have overlapping error bars, the remaining uncertainties are relatively large, illustrating that obtaining highly accurate interaction energies for these large molecules is also challenging for DMC.
As already discussed in Ref. [10], CCSD(T) interaction energies listed Table 2 exhibit large discrepancies compared to DMC for C2C2PD, C3GC and C60@[6]CPPA. In contrast, CCSD(cT)-fit resolves these discrepancies for all systems on the hundred-atom scale, achieving excellent agreement with DMC estimates of Hamdani et. al. [10] to within chemical accuracy () after subtracting the error bars. Even for C60@[6]CPPA, which contains 132 atoms, a discrepancy of only remains, although the error bar of CCSD(cT)-fit is relatively large in this case. We argue that the remaining discrepancies are potentially caused by uncertainties in DMC, CCSD(cT)-fit and the underlying LNO-CCSD(T) calculations. It should be noted that the error bars of LNO-CCSD(T) interaction energies are in some cases underestimated, as exemplified for the Uracil-Uracil PD dimer by the comparison between canonical CCSD(T) and LNO-CCSD(T) interaction energies shown in Fig. 2. Furthermore, the DMC interaction energy of C60@[6]CPPA has not yet been verfied independently using a different DMC implementation as it was done for all other systems listed in Table 2. We also stress that in some cases the differences between the DMC estimates are larger than their respective error bars.
@@ 3 Conclusion
Funding T.S. acknowledges support from the Austrian Science Fund (FWF) [DOI: 10.55776/ESP335]. The computational results presented have been largely achieved using the Vienna Scientific Cluster (VSC). Andreas Irmler and Alejandro Gallo acknowledge support from the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 951786 (The NOMAD CoE). Support and funding from the European Research Council (ERC) (Grant Agreement No. 101087184) is gratefully acknowledged.
Conflict of interest The authors declare no competing interests.
Ethics approval Not applicable
Consent to participate Not applicable
Consent for publication Not applicable
Availability of data and materials The data that support the findings of this study are available from the corresponding authors upon reasonable request.
Code availability The VASP code is a copyrighted software and can be obtained from its official website. The CC4S code is free and open-source and can be downloaded from its website.
Authors’ contributions A.I. and T.S. performed the calculations. A.I. and A.G. implemented the methods in Cc4s and the employed quantum chemistry packages. A.G. advised and implemented related interface methods in the VASP code. All authors wrote the paper.

### v3

@@ 2.2 Estimating the overcorrelation of (T) for weak interactions
In summary, we have demonstrated that CCSD(cT) theory achieves excellent agreement for noncovalent interaction energies between molecular complexes compared to DMC and CCSDT theory. However, we stress that the CC series of methods (CCSD, CCSDT and CCSDTQ) is observed to yield monotonic and rapidly converging interaction energies for small and weakly bound complexes (?). Based on this knowledge, we emphasize that the Q contribution to the interaction energies can be expected to be smaller than its T counterpart, but could possibly yield a significant contribution. Indeed, this is part of the reason for the success of the CCSD(T) approximation for very small molecules, where CCSD(T) is often fortuitously closer to CCSDTQ than CCSDT (?). Here, we argue that this error cancellation no longer functions in the case of large molecular complexes involving strongly polarizable systems such as C2C2PD, C3GC and C60@[6]CPPA. A similar problem is known to occur in MP2 theory, where the truncation of the perturbation series also leads to significantly too strong interaction energies for systems with large polarizability, although MP2 yields relatively accurate interaction energies for systems with an intermediate polarizability (?).
@@
Having demonstrated and explained the reasons for the overestimation of absolute interaction energies on the level of CCSD(T) theory for small molecules with up to 24 atoms and the C2C2PD system, we now want to turn to the discussion of the remaining large molecular complexes where significant absolute discrepancies between DMC and CCSD(T) have been observed. These systems include C3GC from the L7 data set and the C60@[6]CPPA buckyball-ring. Here, substantial differences in the binding energies of and were reported, respectively after subtracting error bars. Although CCSD(cT) calculations for systems of that size are currently not feasible using our approach, we now introduce a simplified model that allows us to estimate the change in interaction energies from CCSD(T) to CCSD(cT) in an approximate manner. Given the linear trend between the different correlation energy contributions to the interaction energy depicted in Fig. 3, it is possible to estimate the (cT) contribution for systems where only MP2, CCSD and (T) are known. These numbers can be calculated using a computationally efficient LNO-CCSD(T) implementation (?). Results computed in this manner are denoted as CCSD(cT)-fit. Details about this procedure and the corresponding error estimates are provided in the supplementary information.
System | CCSD(T) | LNO-CCSD(T) (?) | CCSD(cT) | CCSD(cT)-fit | DMC (?) | DMC (?) |
@@
Table 2 gives our estimated CCSD(cT) interaction energies in comparison to CCSD(T) and DMC findings for seven large molecular complexes. A comparison between CCSD(cT)-fit and the explicitly calculated CCSD(cT) results for GGG, GCGC and C2C2PD shows that the linear regression model is sufficiently reliable for the systems studied in this work. For comparison Table 2 also summarizes the DMC interaction energies from Refs. (?) and (?), which agree to within at least for GGG, C2C2PD, PHE and C3GC. For the remaining systems the DMC estimates show a larger discrepancy and for C60@[6]CPPA only one DMC estimate is available. Although the DMC binding energies have overlapping error bars, the remaining uncertainties are relatively large, illustrating that obtaining highly accurate interaction energies for these large molecules is also challenging for DMC.
As already discussed in Ref. (?), CCSD(T) interaction energies listed Table 2 exhibit large discrepancies compared to DMC for C2C2PD, C3GC and C60@[6]CPPA. In contrast, CCSD(cT)-fit resolves these discrepancies for all systems on the hundred-atom scale, achieving excellent agreement with DMC estimates of Hamdani et. al. (?) to within chemical accuracy () after subtracting the error bars. Even for C60@[6]CPPA, which contains 132 atoms, a discrepancy of only remains, although the error bar of CCSD(cT)-fit is relatively large in this case. We argue that the remaining discrepancies are potentially caused by uncertainties in DMC, CCSD(cT)-fit and the underlying LNO-CCSD(T) calculations. It should be noted that the error bars of LNO-CCSD(T) interaction energies are in some cases underestimated, as exemplified for the Uracil-Uracil PD dimer by the comparison between canonical CCSD(T) and LNO-CCSD(T) interaction energies shown in Fig. 1. Furthermore, the DMC interaction energy of C60@[6]CPPA has not yet been verfied independently using a different DMC implementation as it was done for all other systems listed in Table 2. We also stress that in some cases the differences between the DMC estimates are larger than their respective error bars.
@@ 3 Conclusion
Acknowledgments
Tobias Schäfer acknowledges support from the Austrian Science Fund (FWF) [DOI: 10.55776/ESP335]. The computational results presented have been largely achieved using the Vienna Scientific Cluster (VSC). Andreas Irmler and Alejandro Gallo acknowledge support from the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 951786 (The NOMAD CoE). Support and funding from the European Research Council (ERC) (Grant Agreement No. 101087184) is gratefully acknowledged. We gratefully acknowledge discussions with Adrian Daniel Boese.
Supplementary materials
Materials and Methods
Supplementary Text
Figs. S1 to S2
Tables S1 to S5